# Generalised Bayesian Inference using Robust divergences for von Mises-Fisher distribution

### v1

@@ 7 Conclusion
In this paper, we have developed a robust generalized Bayesian inference for the von Mises-Fisher distribution using the density power divergence and the $\gamma$-divergence. The proposed generalized posterior distributions provide a principled way to incorporate robustness into Bayesian estimation for both the location and concentration parameters, where the presence of outliers can severely distort ordinary Bayesian inference.
@@
Simulation studies across multiple dimensions, sample sizes, and contamination ratios revealed the clear advantage of the proposed methods. While the MSE of the ordinary posterior mean increases sharply as the contamination ratio increases, both the DPD and $\gamma$-D posteriors maintain stable and consistent performance. In particular, the DPD posterior tends to outperform the $\gamma$-D posterior in small-sample and high-dimensional settings, though their performance becomes comparable as the sample size increases. The applications to real wind direction data and the gene expression data also illustrated that the proposed methods yield robust posterior summaries that are not distorted by outlying observations.
Overall, the proposed robust generalized Bayesian inference framework offers a flexible, computationally tractable, and theoretically justified approach for directional data analysis under contamination. Potential extensions of this work include the development of adaptive selection methods for tuning parameters, applications to more complex directional models such as the Fisher-Bingham distribution (Kent,, 1982), and theoretical investigations of robustness in hierarchical or regression-type models (e.g. Holmquist and Gustafsson,, 2017; Scealy and Welsh,, 2011; Scealy and Wood,, 2019) for directional data. These directions represent promising avenues for future research.
Acknowledgment
This work was supported by JSPS KAKENHI Grant Numbers JP23K13019 and JP25H00546.
@@ A.1 Derivative functions and Expectations
We summarize the gradient (score) functions for each loss. The gradient for each loss functions are given by
@@
Also, we summarize the Hessian functions for each loss function, given by
@@ A.3 Selection of tuning parameters
Figure 9: The figures display curve of the asymptotic relative efficiency for the density power-posterior and $\gamma$-posterior means in the case of dimension $p=2$. The dashed red line indicates $95\%$, and the dashed black line indicates the hyperparameter ($\alpha$, $\gamma$) values corresponding to the 95th percentile.

### v2

@@ 7 Conclusion
In this paper, we have developed a robust generalised Bayesian inference for von Mises-Fisher models based on the density power divergence and the $\gamma$-divergence. The proposed generalised posterior distributions provide a principled way to incorporate robustness into Bayesian estimation for both the location and concentration parameters, where the presence of outliers can severely distort ordinary Bayesian inference.
@@
Simulation studies across multiple dimensions, sample sizes, and contamination ratios revealed the clear advantage of the proposed methods. While the MSE of the ordinary posterior mean increases sharply as the contamination ratio increases, both the DPD and $\gamma$-D posteriors maintain stable and consistent performance. In particular, the DPD posterior tends to outperform the $\gamma$-D posterior in small-sample and high-dimensional settings, although their performance becomes comparable as the sample size increases. The applications to real wind direction data and the gene expression data also illustrated that the proposed methods yield robust posterior summaries that are not distorted by outlying observations.
Overall, the proposed robust generalised Bayesian inference framework offers a flexible, computationally tractable, and theoretically justified approach for directional data analysis under contamination. Potential extensions of this work include the development of adaptive selection methods for tuning parameters, applications to more complex directional models such as the Fisher-Bingham distribution (Kent,, 1982), and theoretical investigations of robustness in hierarchical or regression-type models (e.g. Holmquist and Gustafsson,, 2017; Scealy and Welsh,, 2011; Scealy and Wood,, 2019) for directional data. These directions represent promising avenues for future research.
Acknowledgement
This work was supported by JSPS KAKENHI Grant Numbers JP23K13019, JP24K20746, and JP25H00546.
@@ A.1 Derivative functions and Expectations
We summarise the gradient (score) functions for each loss. The gradient for each loss functions are given by
@@
Also, we summarise the Hessian functions for each loss function, given by
@@ A.3 Selection of tuning parameters
Figure 9: The figures display curve of the asymptotic relative efficiency for the density power-posterior and $\gamma$-D posterior means in the case of dimension $p=2$. The dashed red line indicates $95\%$, and the dashed black line indicates the hyperparameter ($\alpha$, $\gamma$) values corresponding to the 95th percentile.